# A Genealogy of Multi-Sensor Foundation Models in Remote Sensing

### v1

@@ 4.3. Different Sensors as Augmentation
Contrastive Radar-Optical Masked Auto-encoders (CROMA) (Fuller et al., 2023) also performs SSL on Sentinel-1 and Sentinel-2 data, training an encoder for each sensor. However, CROMA takes inspiration from the FLIP framework (Jia et al., 2021), which independently masks each mode of data before feeding them into encoders and performing contrastive learning in the standard CLIP (Radford et al., 2021) setup. Masking images before encoding them results in a significantly less memory intensive process than encoding the raw image when vision transformers are used. Therefore, the FLIP framework enables contrastive training with a much larger batch size with the same or less memory footprint, which results in faster convergence. CROMA implements this approach with remotely sensed images, contrastively learning between masked Sentinel-1 and Sentinel-2 images. CROMA also introduces a third encoder that takes in concatenated Sentinel-1 and Sentinel-2 encodings, generates multi-sensor encodings, then is fed into a multi-sensor decoder responsible for constructing the initial Sentinel-1 and Sentinel-2 images. In this manner, CROMA is able to introduce contrastive and reconstruction loss terms into their framework, resulting in a model that effectively learns uni-sensor representations for Sentinel-1 and Sentinel-2 and multi-sensor representations of both Sentinel-1 and Sentinel-2. This framework is benchmarked against BigEarthNet (Sumbul et al., 2019) and FMoW (Christie et al., 2018) classification tasks, demonstrating SOTA performance with images from a single sensor and multi-sensed images. BigEarthNet contains roughly half a million Sentinel-2 images where each patch can have multiple land-cover classes. CROMA also utilizes modified ALiBi positional embeddings (Press et al., 2022) for their ViT encoders; these embeddings provide an advantage over standard sinusoidal positional embeddings by biasing the embeddings based on Euclidean distance between key-query pairs, thus imbuing the embeddings with a degree of spatial awareness.
4.4. Text as Augmentation
CLIP’s original design used text and images as its two modalities, and RemoteCLIP (Liu et al., 2024) goes back to these roots. RemoteCLIP takes in text as its one modality, leaning on image captions to generate natural labels. Other models’ requirements for additional fine-tuning, as opposed to zero-shot utilization, motivates this model’s development, as including text as one of the modalities results in good zero-shot performance downstream (Zhu et al., 2024). This model treats RGB images taken from Sentinel-2, drones, and other aerial platforms as one modality. This variety in the sources for the image modality is designed to result in a more robust encoder that will be resolution agnostic, yet the reported performance takes note of RemoteCLIP’s struggles with zero-shot image classification as an indication that there is additional work to be done in that regard.
For the text modality, RemoteCLIP relies on image captions already present in the human-annotated satellite imagery datasets. Where no captions are paired with an image, a preprocessing stage generates appropriate text annotations. This preprocessing stage performs standard object detection on the image, then uses a script to programmatically generate a corresponding sentence, which is then used as the positive text annotation. Generating captions via script will lack natural variance that is present in a typical corpus, and generating these captions via NLP processes such as ChatGPT will inherently introduce extrinsic noise and potential bias into the dataset. Nevertheless, RemoteCLIP outperforms SOTA models like CLIP and DINOv2 in zero-shot evaluation of object counting in the Remote-Count dataset (Gao et al., 2021), and few-shot classification of 12 RS benchmarks, including EuroSat (Helber et al., 2019), RSC11 (Zhao et al., 2016), and RESISC45 (Cheng et al., 2017), which indicates that descriptive text as a modality has definite value in the remote sensing domain.
@@ 4.4. Text as Augmentation
5. Distillation Networks in Multi-Modal Remote Sensing
@@
One successful example of a distillation network in the RS field is RS-BYOL (Jain et al., 2022b), which borrows BYOL’s framework and uses the multi-modal aspect for multi-sensor remote sensing data. Just as with negative sampling approaches in RS, different sensors’ imaging of the same location are leveraged as natural augmentations of the same image. One modality is drawn from Sentinel-2 optical, whereas the other is drawn from Sentinel-1 SAR. Similarly, additional augmentations of random crops, rotations, color jitter, and grayscale are applied to prevent the model from learning trivial representations of the data.
@@
RS-BYOL utilizes 10 bands of data in Sentinel-2 with 10-20 m resolution and the VV and VH polarizations of SAR present in Sentinel-1. The model outperforms negative sampling approaches such as SeCo in linear probing when all bands are present from Sentinel-2, indicating that there is inherent value in incorporating multispectral bands from Sentinel-2 in addition to the RGB bands. However, SeCo outperforms RS-BYOL with linear probing when only the RGB bands are present, indicating that negative sampling approaches may still have a performance edge in remote sensing. The lighter resource requirements for distillation networks reflects a tradeoff between model performance and hardware requirements.
A sort of Venn diagram where a variety of RS models that utilize contrastive learning are compared with one another to highlight similarities and differences.
6. Masked Auto-Encoders in Remote Sensing
@@
While MAE is a powerful approach in CV, it requires significant adaptations to be effectively leveraged in the RS world. Below are some of the adjustments that have been made to MAE in order to leverage it more effectively in the remote sensing domain (Nedungadi et al., 2024).
@@ 6.1. Masking Approach
The simple CV approach of masking individual patches of an image fails to work effectively with RS imagery (Sun et al., 2023). Compared to natural images in CV, satellite and aerial images in RS often capture small surface features, dense objects, complex backgrounds, and a range of possible object orientations. Similarly, natural images tend to follow expected proportions with well-defined distributions, whereas the same cannot be said for RS images (Rolf et al., 2024). Sun’s Remote Sensing Foundation Model (RingMo) introduces PIMask (Sun et al., 2023) to account for these inherent differences, as masking an entire patch of a RS image might lose information that is unique to that given patch. Therefore, some pixels in each masked patch are randomly preserved, and multilayer convolution is applied to each patch. Compared to traditional CV methods, where masking 50% of an image would completely mask 50% of the patches, masking 50% of an image with PIMask would partially mask a higher percentage of patches. The overall number of pixels masked in the image would still be 50%, but the spatial distribution is different.
@@
MAE in CV finds that masking high percentages of the image (roughly 75%) results in more accurate representations being learned (He et al., 2021), and that carries over to the current implementations of MAE in remote sensing as well (Sun et al., 2023). This parallels with the findings from negative sampling approaches that making the pretext task with contrastive objective more difficult via multiple augmentations results in the model learning better embeddings. RingMo’s embeddings outperform embeddings generated with comparable MAE frameworks and standard CV patch masking, underscoring the importance of improving the masking procedure for satellite imagery.
@@
Spatial-Spectral MAE (S2MAE) (Li et al., 2024a) takes this a step further, as they find that masking 90% of satellite imagery yields the best results in their novel framework. S2MAE trains on multispectral imagery from FMoW (Christie et al., 2018) and BigEarthNet (Sumbul et al., 2019) datasets. Unlike RingMo and other image RGB masking strategies which mask out the same pixel across all channels, S2MAE employs independent 3D masking, which masks out different pixels in each channel. 3D masking was introduced for videos and appears to learn environmental details well, but struggles to track movement from frame to frame (Feichtenhofer et al., 2022). While this is poorly suited for video analysis, it is ideal for RS applications, as seen in S2MAE’s SOTA performance on the Onera Satellite Change Detection (OSCD) (Daudt et al., 2018) dataset and EuroSat (Helber et al., 2019). S2MAE also finds that shallow decoders are poorly suited for SSL with multispectral imagery, a notable departure from established CV practices (He et al., 2021).
@@ 6.2. Temporal Encodings
SatMAE (Cong et al., 2023) leans on MAE’s utilization of a ViT as the backbone, but encodes the hour, month, and year as temporal encodings, then concatenates these transformed temporal encodings to the standard positional encodings that are innate to the functioning of a ViT. As with SeCo, the inclusion of temporal information in self-supervised learning for RS foundation models can have substantial performance improvements, but it also means that the embeddings learned are agnostic to temporal variability, i.e., learning an average of variations over time, thus limiting the value for certain downstream tasks where preservation of temporal nuance is crucial. Still, this ability to encapsulate temporal data highlights the flexibility of the MAE approach in remote sensing.
Taking into account the adjusted positional encoding, SatMAE experiments with different masking approaches. It utilizes consistent masking, where the masked regions are spatially-consistent across all images of a given location, and independent masking, where the location of the patches can change from image to image of a given location. Results indicate that independent masking can trivialize the task, as the model might be able to reconstruct a given masked patch by referring to an unobstructed view of that patch at a given time, so a random cropping augmentation is applied to ensure the model learns robust embeddings.
@@
Prithvi (Jakubik et al., 2023) trains on the Harmonized Landsat Sentinel-2 (HLS) (Claverie et al., 2018) dataset, which contains multispectral images captured by Landsat 8/9 and Sentinel-2 in order to provide a satellite imagery dataset with higher temporal resolution. Prithvi leverages this higher temporal resolution to arrange these 2D images in a time series and updates the traditional MAE positional and patch embeddings to be 3-dimensional, where the third dimension pertains to time. When masking patches, they specifically apply 3D convolutions to avoid losing information specific to any one given patch, similar to RingMo’s (Sun et al., 2023) approach. Aside from the incorporation of temporal data, Prithvi also utilizes a landscape stratified sampling strategy to avoid bias towards more common ecosystems or landscapes.
@@
Presto (Tseng et al., 2024) takes this approach even further, constructing a pixel-timeseries training dataset that consists of input from multiple types of data, including Sentinel-2 RGB, Sentinel-1 SAR, DEM, and the Dynamic World (Brown et al., 2022) dataset, which contains land cover labels for 10m resolution Sentinel-2 imagery. The framework then employs a variety of different strategies for training, masking channels, blocks of timesteps, single timesteps, and/or random pixels for reconstruction by a downstream encoder. The resulting encoder is performant with regards to feature extraction, fine-tuning, and time series forecasting. The variety of masking strategies employed ensures that Presto can successfully process time series datasets with missing channels, a variety of temporal resolutions, and only a small subset of timesteps. However, performance on tasks that are temporally agnostic and rely on a single image lags behind SOTA models; the authors note that this is possibly due to the model’s inability to process images with a spatial resolution higher than 10m.
6.3. Scale
@@
Foundation models in CV, and most foundation models in RS by extension, process images in local image coordinates, meaning that the learned representations are dependent on the scale and resolution of the satellite sensor. ScaleMAE (Reed et al., 2023) is designed to address this issue by scaling the image coordinate system such that it matches the actual spatial scale of objects, ensuring that patch size, as measured in geospatial units, is the same across various images. This is done via introducing a Ground Sample Distance (GSD) positional encoding, which then informs the ViT of both the geospatial scale and image position of the patch. Put differently, ScaleMAE addresses relative spatial scale, but does not incorporate absolute geolocation in its training. As with SatMAE, ViT’s framework, with dual patch and positional encodings, lends itself well to introducing the concept of scale to the model.
Dovetailing with the incorporation of scale, ScaleMAE’s decoder produces both high and low resolution images in order to capture both low-frequency features, such as mountains and rivers, and more fine-grained high-frequency features, such as vegetation. It does so by utilizing a Laplacian-pyramid decoder instead of the standard transformer decoder.

### v2

@@ 5.3. Multisensor Pairing
5.4. Image-Text Pairing
CLIP’s original design used text and images as its two modalities, and RemoteCLIP (Liu et al., 2024) goes back to these roots. RemoteCLIP takes in text as its one modality, leaning on image captions to generate natural labels. Other models’ requirements for additional fine-tuning, as opposed to zero-shot utilization, motivates this model’s development. Including text as one of the modalities results in good zero-shot performance downstream (Zhu et al., 2024). This model treats RGB images taken from Sentinel-2, drones, and other aerial platforms as one modality. This variety in the sources for the image modality is designed to result in a more robust encoder that will be resolution agnostic. RemoteCLIP’s reported underperformance in zero-shot image classification tasks indicates that there is additional work to be done in that regard.
For the text modality, RemoteCLIP relies on image captions already present in the human-annotated satellite imagery datasets. Where no captions are paired with an image, a preprocessing stage generates appropriate text annotations. This preprocessing stage performs standard object detection on the image, then uses a script to programmatically generate a corresponding sentence. That sentence is then used as the positive text annotation. Generating captions via script will lack natural variance that is present in a typical corpus, and generating these captions via NLP processes such as ChatGPT will inherently introduce extrinsic noise and potential bias into the dataset. Nevertheless, RemoteCLIP outperforms SOTA models like CLIP and DINOv2 in zero-shot evaluation of object counting in the Remote-Count dataset (Gao et al., 2021), and few-shot classification of 12 RS benchmarks, including EuroSat (Helber et al., 2019), RSC11 (Zhao et al., 2016), and RESISC45 (Cheng et al., 2017). This performance indicates that text as a modality has definite value in the remote sensing domain.
@@ 5.4. Image-Text Pairing
6. Distillation Networks in Remote Sensing
Given the requirement for handling a sizable amount of remotely sensed data, efficiency-oriented SSL approaches like distillation have become popular for remote sensing foundation models.
6.1. Sensor-Agnostic Encoder
@@
One successful example of a distillation network in the RS field is RS-BYOL (Jain et al., 2022b), which borrows BYOL’s framework and uses the multi-modal aspect for multi-sensor remote sensing data. Just as with negative sampling approaches in RS, different sensors’ imaging of the same location are leveraged as natural augmentations of the same image. One modality is drawn from Sentinel-2 optical, whereas the other is drawn from Sentinel-1 SAR. Similarly, additional augmentations of random crops, rotations, color jitter, and grayscale are applied to prevent the model from learning trivial representations of the data. RS-BYOL utilizes 10 bands of data in Sentinel-2 with 10-20 m resolution and the VV and VH polarizations of SAR present in Sentinel-1. RS-BOYL only trains a single encoder at a time and chooses 3 random bands of Sentinel-2 and 3 random polarizations from Sentinel-1 to make up the two image views. They perform distillation with view 1 passing through the student encoder and view 2 passing through the teacher encoder, then also perform distillation with the views passing through the opposite encoders. The resulting framework represents an early attempt to learn sensor-invariant features.
@@
The model outperforms negative sampling approaches such as SeCo in linear probing when all bands are present from Sentinel-2, indicating that there is inherent value in incorporating multispectral bands from Sentinel-2 in addition to the RGB bands. However, SeCo outperforms RS-BYOL with linear probing when only the RGB bands are present, indicating that negative sampling approaches may still have a performance edge in remote sensing. The lighter resource requirements for distillation networks reflects a tradeoff between model performance and hardware requirements. RS-BYOL represents a somewhat straightforward adaptation of CV distillation SSL techniques in RS, but other approaches make more dramatic changes.
DINO-MM (Wang et al., 2022a) is another RS foundation model leveraging distillation as its SSL objective, although this model is built off of DINO’s (Caron et al., 2021) framework instead. DINO-MM trains on BigEarthNet-MM (Sumbul et al., 2019), a dataset containing roughly half a million pairs of Sentinel-2 and Sentinel-1 images at the same spatiotemporal scenes. DINO-MM represents one of the first RS foundation models to explore the use of ViTs (Dosovitskiy et al., 2021) as a backbone. As such, its student and teacher encoders can handle flexible input shapes. DINO-MM takes full advantage of this, stacking SAR and MS channels on top of one another to create the ’original image’. DINO-MM applies the standard augmentations present in DINO, but also introduces a RandomSensorDrop augmentation that results in one of three outcomes:
All SAR channels are dropped
All MS channels are dropped
No channels are dropped
This results in an encoder that is capable of leveraging whatever satellite imagery is available when creating embeddings. In theory, it also means that this model learns intra- and inter-modal representations of the data. However, this framework’s flexibility does not necessarily guarantee performance when all modes of data are present. Reported ablation studies show that downstream assessment on BigEarthNet’s classification task perform well with Sentinel-2 data, but do not actually improve when given additional Sentinel-1 information. In the case of top-1 benchmarking, the model’s performance actually decreased slightly while fine-tuning on Sentinel-1 and Sentinel-2 imagery (when compared to purely Sentinel-2 fine-tuning). It is possible that benchmarking on tasks where SAR data is more relevant, such as Sen1Floods11 (Bonafilia et al., 2020), might yield more favorable results with multiple modes of data present.
6.2. Multiple Contrastive Tasks
SkySense (Guo et al., 2024) represents a significant step in RS foundation model size, with 2.09 billion trainable parameters. Aside from model size, SkySense performs distillation SSL at multiple stages of its training pipeline, resulting in a model capable of learning effective high level and low level features. SkySense trains on three modalities: time series of Sentinel-1, time series of Sentinel-2, and high resolution RGB imagery from Worldview 3 and 4. For each mode, there is a teacher and student encoder. SkySense performs two sets of random augmentations per image, then sends the different views to the student and teacher. It performs multi-modal contrastive loss between the embeddings from the student encoders and then performs a number of contrastive learning tasks between student and teacher embeddings, namely:
pixel-level contrastive loss
image-level contrastive loss via average pooling of pixels
object-level contrastive loss by using the Sinkhorn-Knopp algorithm (Knight, 2008) to perform unspervised clustering on pixel-level features
After these contrastive tasks, the embeddings go through a multi-modal temporal fusion encoder at both the student and teacher level. This results in a fused embedding. Sinkhorn-Knopp is applied again, but this time, the resulting clusters are compared to geographic regions in order to learn geo-context. Then the fused embeddings go through the same contrastive learning tasks described in the paragraph above. The resulting model was reported to achieve SOTA performance on downstream tasks.
7. Redundancy Reduction in Remote Sensing
Decoupling Common and Unique Representations (DeCUR) for Multimodal SSL (Wang et al., 2024) builds on the Barlow Twins redundancy reduction approach, but modifies it to handle multiple modes of data. DeCUR explores several multi-modal training datasets. In order to explore the efficacy of their multi-modal training approach, DeCUR trains one model per each of the following datasets:
DeCUR’s key distinction from other multi-modal approaches we have discussed lies in its desire to learn inter-modal representations, but preserve data unique to a given modality. It does so by setting aside a percentage of the latent space for shared representations and leaving the other percentage of the latent space for modality specific information. When the Barlow Twin’s redundancy reduction loss is applied from one mode to another, DeCUR only calculates this loss for shared dimensions of the latent space. DeCUR also calculates redundancy reduction loss within a given mode itself, and does so for all of the dimensions within the latent space. This strategy enables DeCUR to align modalities without losing modality-specific information and results in SOTA performance on downstream tasks.
One potential downside of this approach is the calculation of the shared latent space dimensions. The authors use a light grid search in each dataset to determine how much of the latent space should be used for learning shared representations. This percentage tends to hover around 75-80% in the datasets above, but calculating it becomes costly with larger datasets. The authors also found no significant drop in performance when they dropped the shared dimensions to as low as 50%. This suggests a degree of sparsity in the shared latent space and an opportunity to refine how the shared dimension percentage is calculated, or ways to even better leverage all dimensions.
8. Masked Image Modeling in Remote Sensing
@@
While MIM is a powerful approach in CV, there is significant room to optimize performance for remotely sensed imagery. Below are some of the adjustments that have been made to MIM in order to leverage it more effectively in the remote sensing domain. Most models discussed follow the MAE masking and decoder paradigm, with a few noted exceptions.
@@ 8.1. Masking Approach
The simple CV approach of masking individual patches of an image can work with RS imagery, but there is room for improvement (Sun et al., 2023). Compared to natural images in CV, satellite and aerial images in RS often capture small surface features, dense objects, complex backgrounds, and a range of possible object orientations. Similarly, natural images tend to follow expected proportions with well-defined distributions, whereas the same cannot be said for RS images (Rolf et al., 2024). Sun’s Remote Sensing Foundation Model (RingMo) introduces PIMask (Sun et al., 2023) to account for these inherent differences, as masking an entire patch of a RS image might lose information that is unique to that given patch. Therefore, some pixels in each masked patch are randomly preserved, and multilayer convolution is applied to each patch. Compared to traditional CV methods, where masking 50% of an image would completely mask 50% of the patches, masking 50% of an image with PIMask would partially mask a higher percentage of patches. The overall number of pixels masked in the image would still be 50%, but the spatial distribution is different. RingMo is officially classified as a MIM approach, as the masking strategy does not allow for the encoder to only be fed unmasked patches.
@@
MAE in CV finds that masking high percentages of the image (roughly 75%) results in more accurate representations being learned (He et al., 2021), and that carries over to the current implementations of MAE in remote sensing as well (Sun et al., 2023). This parallels with the findings from negative sampling approaches that making the pretext task with contrastive objective more difficult via multiple augmentations results in the model learning better embeddings. RingMo’s embeddings outperform embeddings generated with comparable MIM frameworks and standard CV patch masking, underscoring the importance of improving the masking procedure for satellite imagery.
@@
Spatial-Spectral MAE (S2MAE) (Li et al., 2024a) takes this a step further, as they find that masking 90% of satellite imagery yields the best results in their novel framework. S2MAE trains on multispectral imagery from FMoW (Christie et al., 2018) and BigEarthNet (Sumbul et al., 2019) datasets. Unlike RingMo and other image RGB masking strategies which mask out the same pixel across all channels, S2MAE employs independent 3D masking, which masks out different pixels in each channel. 3D masking was introduced for videos and appears to learn environmental details well, but struggles to track movement from frame to frame (Feichtenhofer et al., 2022). While this is poorly suited for video analysis, it is ideal for RS applications, as seen in S2MAE’s SOTA performance on the Onera Satellite Change Detection (OSCD) (Daudt et al., 2018) and EuroSat (Helber et al., 2019) benchmarks. S2MAE also finds that shallow decoders are poorly suited for MAE with multispectral imagery, a notable departure from established CV practices (He et al., 2021).
SatMAE (Cong et al., 2023) also experiments with different masking approaches. It utilizes consistent masking, where the masked regions are spatially-consistent across all images of a given location, and independent masking, where the location of the patches can change from image to image of a given location. Results indicate that independent masking can trivialize the task by leveraging spectral redundancy. In other words, the model might be able to reconstruct a given masked patch by referring to an unobstructed view of that patch at a given time. A random cropping augmentation is applied to address this and ensure the model learns robust embeddings.
SpectralGPT (Hong et al., 2024) experiments with a unique 3D-cube masking strategy. Training on 12 bands of Sentinel-2 data from FMoW (Christie et al., 2018) and BigEarthNet (Sumbul et al., 2019), SpectralGPT’s approach cannot be described as either independent or consistent masking. Instead of tokenizing channel by channel, SpectralGPT creates non-overlapping tokens which span 3 channels. Independent masking is then applied to each layer of tokens. This results in a blend of consistent and independent masking that is informed by spectral resolution. SpectralGPT also uses multiple loss terms in their approach, generating loss based on token-to-token reconstruction, but also spectral signature-to-spectral signature reconstruction. Similar to S2MAE (Li et al., 2024a), SpectralGPT finds that it performs best with a masking rate of 90% and a deeper decoder.
@@ 8.2. Temporal Encodings
SatMAE (Cong et al., 2023) leans on MAE’s utilization of a ViT as the backbone, but encodes the hour, month, and year as temporal encodings. SatMAE then concatenates these temporal encodings to the standard positional encodings that are innate to the functioning of a ViT. As with SeCo, the inclusion of temporal information in self-supervised learning for RS foundation models can have substantial performance improvements, but it also means that the embeddings learned are agnostic to temporal variability. Thus, the embeddings’ applicability to temporally sensitive downstream tasks is limited. Still, this ability to encapsulate temporal data highlights the flexibility of the MAE approach in remote sensing.
@@
Prithvi (Jakubik et al., 2023) trains on the Harmonized Landsat Sentinel-2 (HLS) (Claverie et al., 2018) dataset, which contains multispectral images captured by Landsat 8/9 and Sentinel-2. Given the differing revisit times of the two satellites, HLS is a satellite imagery dataset with a high temporal resolution. Prithvi leverages this higher temporal resolution to arrange these 2D images in a time series and updates the traditional MAE positional and patch embeddings to be 3-dimensional, where the third dimension pertains to time. When masking patches, they specifically apply 3D convolutions to avoid losing information specific to any one given patch, similar to RingMo’s (Sun et al., 2023) approach. Unlike RingMo, Prithvi fully masks patches. Prithvi also utilizes a landscape stratified sampling strategy to avoid bias towards more common ecosystems or landscapes.
@@
Presto (Tseng et al., 2024) takes this approach even further, constructing a pixel-timeseries training dataset that consists of input from multiple types of data, including Sentinel-2 RGB, Sentinel-1 SAR, DEM, and the Dynamic World (Brown et al., 2022) dataset. Dynamic World contains land cover labels for 10m resolution Sentinel-2 imagery. The framework then employs a variety of different strategies for training. It masks channels, blocks of timesteps, single timesteps, and/or random pixels for reconstruction by a downstream encoder. The resulting encoder performs well on feature extraction, fine-tuning, and time series forecasting. The variety of masking strategies ensures that Presto can successfully process time series datasets with missing channels, a variety of temporal resolutions, and only a small subset of timesteps. However, performance on tasks that are temporally agnostic and rely on a single image lags behind SOTA models. The authors note that this is possibly due to the model’s inability to process images with a spatial resolution higher than 10m.
8.3. Spatial Resolution
@@
Foundation models in CV, and most foundation models in RS by extension, process images in local image coordinates. This means that the learned representations are dependent on the spatial resolution of the satellite sensor. ScaleMAE (Reed et al., 2023) is designed to address this issue by scaling the image coordinate system such that it matches the actual spatial scale of objects, ensuring that patch size, as measured in geospatial units, is the same across various images. This is done via introducing a Ground Sample Distance (GSD) positional encoding, which then informs the ViT of both the geospatial scale and image position of the patch. Put differently, ScaleMAE addresses relative spatial resolution, but does not incorporate absolute geolocation in its training. As with SatMAE, ViT’s framework lends itself well to introducing the concept of spatial resolution to the model. Dovetailing with the incorporation of spatial resolution, ScaleMAE’s decoder produces both high and low resolution images in order to capture both low-frequency features, such as mountains and rivers, and more fine-grained high-frequency features, such as vegetation. It does so by utilizing a Laplacian-pyramid decoder (Burt and Adelson, 1983) instead of the standard transformer decoder.
Where ScaleMAE focuses on understanding spatial resolution of the overall image, USat (Irvin et al., 2023) chooses to focus on spatial resolution across image channels. Different bands of the same sensor can have dramatically different spatial resolutions. For instance, in the case of Sentinel-2, bands (i.e., input channels) range from 10m resolution to 60m resolution. Many approaches up- or down- sample channels to yield a consistent spatial resolution, and some simply drop low resolution channels altogether (Jain et al., 2022a). This ensures that patch sizes remain consistent, but may lose resolution specific information. USat trains on Sentinel-2 and NAIP, and adjusts the number of patches per channel depending on the spatial resolution. High resolution channels have more patches to capture specific high resolution information. Lower resolution channels have fewer patches, but the total number of patches must be evenly divisible by the number of high resolution patches. This enables channels with different resolutions to cleanly map to one another. Each channel will be masked by the same percentage. More patches are masked in high resolution channels to preserve the percentage of the image being masked. Then, patches are aggregated via spectral pooling, and a standard decoder is used to reconstruct the original image. This approach effectively extracts high-level and low-level features from differing resolutions, but does have some drawbacks. Namely, the assumption that different sensors will have a resolution that is cleanly divisible will not always be true. The authors also note that the approach becomes costly to train with multiple different sensors.
SatMAE++ (Noman et al., 2024) takes inspiration from ScaleMAE’s approach and seeks to introduce spatial resolution awareness to models that train on non-RGB modalities, training on Sentinel-2 imagery in FMoW (Christie et al., 2018). They go about doing so with a more straightforward approach than ScaleMAE, doing away with GSD encodings and Laplacian Pyramid decoders. Instead, SatMAE++ has a multi-resolution reconstruction objective. An image is upsampled 2x and 4x, masking occurs, and then the same decoder is responsible for reconstructing different resolution images based on the corresponding masked input. For the original image, standard mean squared error is applied. For the upsampled blocks, L1 loss is used.